1Find the vertex and focus of the parabola whose equation is 4y = x2 + 4.
aV(0, 4), F(0, 3)
bV(0, 1), F(0, 2)
cV(4, 0), F(3, 0)
dV(1, 0), F(2, 0)

2Find the center and radius of the circle whose equation is x2 + 10x + y2 = 75.
aC(–10, 0), r = 100
bC(–10, 0), r = 10
cC(–5, 0), r = 100
dC(–5, 0), r = 10

3Find the foci of the ellipse with the following equation.
((x-1)ˆ2÷9)+((Y+2)ˆ2÷25)=1
aF1(5, –2), F2(–3, –2)
bF1(1, 2), F2(1, –6)
cF1(4, –2), F2(–2, –2)
dF1(1, 1), F2(1, –5)

4Find the slopes of the asymptotes of a hyperbola with the following equation.
(Yˆ2÷81)-(xˆ2÷64)=1
a8/9
b9/8
c8/9, –8/9
d9/8, –9/8

Answer by lwsshak3(6494)

1Find the vertex and focus of the parabola whose equation is 4y = x2 + 4.
2Find the center and radius of the circle whose equation is x2 + 10x + y2 = 75.
3Find the foci of the ellipse with the following equation.
((x-1)ˆ2÷9)+((Y+2)ˆ2÷25)=1
4Find the slopes of the asymptotes of a hyperbola with the following equation.
(Yˆ2÷81)-(xˆ2÷64)=1
***
1) Find the vertex and focus of the parabola whose equation is 4y = x^2 + 4.
y=x^2/4+4/4
y=x^2/4+1
y=(1/4)(x^2)+1
x^2/4=y-1
x^2=4(y-1)
This is an equation of a parabola of standard form: (x-h)^2=4p(y-k), (h, k) being the (x, y) coordinates of the vertex. Parabola has a vertical axis of symmetry and it opens upwards.
For given equation:
vertex: (0,1)
..
4p=4
p=1
focus: (0,2)
ans:bV(0, 1), F(0, 2)
..
2) Find the center and radius of the circle whose equation is x2 + 10x + y2 = 75.
x^2+10x+y^2=75
complete the square
(x^2+10x+25)+(y^2)=75+25=100
(x+5)^2+(y+0)^2=100
center:(-5,0)
radius=√100=10
ans:dC(–5, 0), r = 10
..
3) Find the foci of the ellipse with the following equation.
((x-1)ˆ2/9)+((Y+2)ˆ2/25)=1
This is an equation of an ellipse with vertical major axis of the standard form:
(x-h)^2/b^2+(y-k)^2/a^2=1, a>b, with (h,k) being the (x,y) coordinates of the center.
center: (1,-2)
a^2=25
b^2=9
c^2=a^2-b^2=25-9=16
c=√16=4
Foci: (1,-2ħc)=(1,-2ħ4)=(1,2) and (1,-6)
ans: bF1(1, 2), F2(1, –6)
..
4Find the slopes of the asymptotes of a hyperbola with the following equation.
(Yˆ2/81)-(xˆ2/64)=1
This is an equation of a hyperbola with vertical transverse axis of the standard form:
(y-k)^2/a^2-(x-h)^2/b^2=1, with (h,k) being the (x,y) coordinates of the center
For given equation:
Center: (0,0)
a^2=81
a=√81=9
..
b^2=64
b=√64=8
slope of asymptotes=ħa/b=ħ9/8
ans:
d9/8, –9/8